A bird flying due north with a velocity at 15m/s shed a feather in air current flowing at 20m/s due east. Find the resultant and direction of the feather.


1
Expert's answer
2022-01-18T09:36:23-0500

"v=\sqrt{v_1^2+v_2^2}=25~\frac ms,"

"\alpha=\arctan\frac{v_2}{v_1}=53°" E to N.